To apply Cramer's rule to solve the following system of equations:
2x₁+x₂+x₃=4
x₁-x₂+2x₃=2
3x₁-2x₂-x₃=0
1
Expert's answer
2019-07-09T11:40:27-0400

d=2*((-1)*(-1)-(-2)*2)-1*(1*(-1)-(-2)*1)+3*(1*2-(-1)*1) = 18 

d1=4*((-1)*(-1)-(-2)*2)-2*(1*(-1)-(-2)*1)+0*(1*2-(-1)*1) = 18 

d2=2*(2*(-1)-0*2)-1*(4*(-1)-0*1)+3*(4*2-2*1) = 18 

d3=2*(2*(-1)-0*2)-1*(4*(-1)-0*1)+3*(4*2-2*1) = 18 

x1=d1/d=1

x2=d2/d=1

x3=d3/d=1